OpenStudy (anonymous):

Use properties of logarithms to expand the logarithmic expression as much as possible. Where possible, evaluate logarithmic expressions without using a calculator.

OpenStudy (anonymous):

\[\ln \sqrt[5]{ey}\]

OpenStudy (anonymous):

The exponent its a 5

OpenStudy (inkyvoyd):

ln(ab)=ln a+ln b

OpenStudy (anonymous):

ln(a^b)=blna use this first

OpenStudy (inkyvoyd):

remember

OpenStudy (inkyvoyd):

the fifth root is equal to ^(1/5)

OpenStudy (anonymous):

my knee?

OpenStudy (inkyvoyd):

lol

OpenStudy (anonymous):

ohh.

OpenStudy (anonymous):

Dont get it

OpenStudy (inkyvoyd):